OpenStudy (anonymous):

e started a exercise program in which each week he works out at the gym between 4.5 and 6 hours and walks between 9 and 12 miles. 1)Write the system of inequalites and graph to show the number of hours diego works out and walks per week

OpenStudy (anonymous):

can you help me please

OpenStudy (mertsj):

\[4.5\le g \le 6\] \[9\le w \le 12\]